OpenStudy (anonymous):

Which is 1024 written as a power of a number? A.54 B.2564 C.5122 D.45

OpenStudy (anonymous):

1024 is \(2^{10}\) which is the same as \(\left(2^2 \right)^5 = 4^5\) So D